Determine the numerical value of Keq at 25 °C for the electrochemical reactions below. I haven’t lectured on this in class, but the “how-to” is clearly illustrated in the textbook in Chapter 19. (a) PbO2(s) + 4 H+(aq) + 2 Cl–(aq) → ← Pb2+(aq) + 2 H2O(l) + Cl2(g) (acidic solution) Keq____________________________ (b) 3 O2(g) + 2 Br–(aq) → ← 2 BrO3–(aq) (basic solution) Keq____________________________
Answer #1

Answer -

Temperature = 25\degreeC

Keq = ?

Some Useful information,

Species \DeltaG\degreef (kJ/mol)
PbO2(s) -215.476
H+(aq) 0
Cl(aq) -131.25208
Pb2+(aq) -24.43
H2O(l) -237.178408
Cl2(g) 0
O2(g) 0
Br(aq) -103.9724
BrO3(aq) 1.6736

a)

PbO2(s) + 4 H+(aq) + 2 Cl(aq) \rightleftharpoons Pb2+(aq) + 2 H2O(l) + Cl2(g)

We know that,

\DeltaG\degree = -RTlnKeq

Keq = e-\DeltaG\degree/RT ---------------------A

where,\DeltaG\degree = Gibbs Free Energy

R = 8.314 Jmol-1K-1

T = Temperature

PbO2(s) + 4 H+(aq) + 2 Cl(aq) \rightleftharpoons Pb2+(aq) + 2 H2O(l) + Cl2(g)

Also,

\DeltaG\degreerxn = \sum n * \Delta G\degreef(products) - \sum n * \Delta G\degreef(reactants)

where, n = stiochiometric coefficient

\DeltaG\degreerxn = 1 * \Delta G\degreef(Pb2+(aq)) + 2 * \Delta G\degreef(H2O(l)) + 1 *\DeltaG\degreef(Cl2(g)) - 1* \Delta G\degreef(PbO2(s)) - 4 *\DeltaG\degreef(H+(aq)) - 2* \Delta G\degreef(Cl(aq))

\DeltaG\degreerxn = [1 * (-24.43 kJ/mol)] + [2 * (-237.129 kJ/mol)] + [1 * (0 kJ/mol)] - [1* (-215.476 kJ/mol)] - [4 * (0 kJ/mol)] - [2 * (-131.25208 kJ/mol)]

\DeltaG\degreerxn = -20.70784 kJ/mol or -20707.84 J/mol [1kJ/mol = 1000J/mol]

Put this in A,

Keq = e-(-20707.84 J/mol/(8.314 Jmol-1K-1 * 298K)

Keq = e8.36

Keq = 4272.69 or 4.27 * 10-3 [Answer]

b)

3 O2(g) + 2 Br(aq)  \rightleftharpoons 2 BrO3(aq)

We know that,

\DeltaG\degree = -RTlnKeq

Keq = e-\DeltaG\degree/RT ---------------------A

where,\DeltaG\degree = Gibbs Free Energy

R = 8.314 Jmol-1K-1

T = Temperature

3 O2(g) + 2 Br(aq)  \rightleftharpoons 2 BrO3(aq)

Also,

\DeltaG\degreerxn = \sum n * \Delta G\degreef(products) - \sum n * \Delta G\degreef(reactants)

where, n = stiochiometric coefficient

\DeltaG\degreerxn = 2 * \Delta G\degreef(BrO3(aq)) - 3* \Delta G\degreef(O2(g)) - 2 *\DeltaG\degreef(Br(aq))

\DeltaG\degreerxn = [2 * (1.6736 kJ/mol)] - [3 * (0 kJ/mol)] - [2 * (-103.9724 kJ/mol)]

\DeltaG\degreerxn = 211.292 kJ/mol or 211292 J/mol [1kJ/mol = 1000J/mol]

Put this in A,

Keq = e-(211292 J/mol/(8.314 Jmol-1K-1 * 298K)

Keq = e85.28

Keq = 1.09 * 1037 [Answer]